How do I evaluate this equation? -50+(-7/9)+35+(-11/9)=
Answer by tommyt3rd(5050)

We actually evaluate expressions and solve equations (in case you need to know for future reference). So we are just writing what simpler quantities have the same value as our expression until we cannot find a simpler one.
-50+(-7/9)+35+(-11/9)=
-50+35 -7/9 -11/9 =
-15 - 18/9 =
-15 - 2 =
-17